-4y+10=-2(3y+1 I need to solve for x and I dont understand how
Found 2 solutions by papugai, Alan3354:
Answer by papugai(18)

-4y+10=-2(3y+1)
Multiply each term in the bracket by -2:
-4y+10=-2*3y+(-2)*1
-4y+10=-6y-2
Add 6y to both sides:
2y+10=-2
Subtract 10 from both sides:
2y=-12
Divide both sides by 2:
y=-6

Answer by Alan3354(69443)

There's no x in the equation.
If you meant y, someone has done that.